40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 18 10:46:51 2010, Koji, Update, ABSL, PLL reconstructed 
    Reply  Thu Mar 18 12:27:56 2010, Alberto, Update, ABSL, PLL reconstructed 
       Reply  Thu Mar 18 13:40:35 2010, Koji, Update, ABSL, PLL reconstructed 
          Reply  Thu Mar 18 21:42:26 2010, Koji, Update, ABSL, PLL reconstructed 
             Reply  Mon Mar 22 22:11:26 2010, Mott, Update, ABSL, PLL reconstructed IMG_0150.JPG
                Reply  Mon Mar 22 23:37:32 2010, Mott, Update, ABSL, PLL reconstructed 2010-03-22_23.14.00.jpg2010-03-22_23.24.50.jpg
                   Reply  Tue Mar 23 18:44:46 2010, Mott, Update, ABSL, PLL reconstructed NPRO_PLL_freqresp.png
                      Reply  Wed Mar 24 03:58:18 2010, kiwamu, matt, koji, Update, Green Locking, leave PLL locked DSC_1402.JPG
                         Reply  Wed Mar 24 04:22:51 2010, kiwamu, matt, Update, Green Locking, two NPRO PLL  
                            Reply  Wed Mar 24 14:52:02 2010, Alberto, Update, Green Locking, two NPRO PLL  
                         Reply  Wed Mar 24 15:59:59 2010, kiwamu, matt, Update, Green Locking, leave PLL locked 
                            Reply  Thu Mar 25 17:29:48 2010, kiwamu, mott, Update, Green Locking, PLL two NPROs 
Message ID: 2697     Entry time: Mon Mar 22 23:37:32 2010     In reply to: 2696     Reply to this: 2703
Author: Mott 
Type: Update 
Category: ABSL 
Subject: PLL reconstructed 

 

We have managed to lock the PLL to reasonable stability. The RF input is attenuated by 26 dBm and the beat signal locks very close to the carrier of the marconi (the steps on the markers of the spectrum analyzer are coarse).  We can use the marconi and the local boost of the pdh box to catch the lock at 0 gain.  Once the lock is on, the gain can be increased to stabilize the lock.  The locked signals are shown in the first photo (the yellow is the output of the mixer and the blue is the output to the fast input of the laser.  If the gain is increased too high, the error signal enters an oscillatory regime, which probably indicates we are overloading the piezo.  This is shown in the second photo, the gain is being increased in time and we enter the non-constant regime around mid-way through.

Tomorrow I will use this locked system to measure the PZT response (finally!).

Attachment 1: 2010-03-22_23.14.00.jpg  1.050 MB  | Show | Show all
Attachment 2: 2010-03-22_23.24.50.jpg  1.060 MB  | Show | Show all
ELOG V3.1.3-